OpenStudy (anonymous):

The y-intercept of 2x + 3y = 15 is (0, ___).

OpenStudy (agreene):

solve for y. 2x+3y=15 3y=15-2x divide by 3 y= (15-2x)/3 put 0 in for x y= (15-2(0))/3 = 5 intercept is 5 QED